# Investigating LLMs in Clinical Triage: Promising Capabilities, Persistent Intersectional Biases

### v1

@@ 4 Results
Additionally, the smaller Llama-3.1-8B-Instruct model demonstrates competitive performance. While its performance is limited on the ESI-Handbook dataset due to the scarcity of training samples, it performs well on KTAS and MIMIC, where larger training sets are available, ranking third-best in terms of QWK and MSE.
@@
MIMIC. The MIMIC dataset is a rich resource with hundreds of thousands of samples, providing the human triage distribution at scale. One striking observation, illustrated in Figure 2, is the discrepancy between human-assigned triage labels in MIMIC and the predictions made by LLMs. Specifically, LLMs exhibit hesitation in assigning patients to the most severe acuity level (1), a pattern previously documented in other tasks (Li et al. 2024b). This discrepancy underscores a broader issue of calibration.
However, as noted earlier, a critical distinction between MIMIC and the other datasets is the quality of the labels themselves. Unlike KTAS, where triage labels are established through expert consensus, or the ESI Handbook, where labels are curated with definitive ground truth, MIMIC labels are derived from individual human decisions. This may partially explain why CoT and AutoCoT, which improves performance on KTAS and ESI, degrades performance on MIMIC since the labels themselves are faulty. Further compounding this issue, MIMIC provides less clinical information than KTAS (which includes a nurse’s diagnosis) or the ESI Handbook (which presents structured scenarios).
@@
Intersectional Biases. Our counterfactual analysis reveals significant demographic preferences across all evaluated models, with consistent patterns emerging across both sex and racial groups. Notably, these biases manifest more strongly at certain intersections of sex and race. For instance, we find that Native Hawaiian women are significantly favored (lower triage scores are prioritized) by GPT and Claude models. In the case of GPT-4o-mini with vanilla prompting, the disparity is particularly stark, with an average acuity level of 2.38 for Native Hawaiian women compared to 3.05 for white men—a nearly full triage level difference.
On that note, interestingly, White men emerge as the least preferred demographic across all models—while this effect is not as extreme as other biases, it is notably consistent. Unlike other groups that may be randomly favored by specific models, white men are never the most preferred group. Additionally, Claude models exhibit a broader preference for Native Hawaiian, Hispanic, and Black patients across both sexes. Some biases appear model-specific such as GPT-4o which favors Hispanic men unlike its mini counterpart.
@@
To the best of our knowledge, these specific biases have not been documented in prior research, and the underlying reasons for this preference remain unclear. However, the statistical significance of these findings highlights the need for further investigation. Given that these biases contradict societal disparities typically observed in healthcare, we hypothesize that they may be a downstream effect of post-training alignment. Interestingly, we observe a pattern among models that exhibit weaker or no biases. CoT prompting consistently reduces the bias, suggesting that explicit reasoning enables models to overlook irrelevant demographic attributes. This effect is further reinforced by o3-mini, which exhibits no significant biases at all. Moreover, Gemini-2.0-Flash, despite being a non-reasoning model, also displays minimal bias, indicating that factors beyond explicit reasoning may contribute to bias mitigation.
Ablations
@@
KTAS MIMIC
Serialization F-1 Score F-1 Score
Commas 0.569 0.582
Natural 0.511 0.585
Newlines 0.499 0.580
Spaces 0.540 0.582
Table 3: Evaluation of Serialization Strategies on KTAS and MIMIC.
@@ 5 Discussion
Our findings highlight the emergence of demographic patterns in LLM predictions. A straightforward mitigation strategy could involve masking sex and race attributes from model inputs. In the case of triage, where demographic factors should not influence prioritization, removing them appears to be a reasonable approach. However, this strategy is context-dependent, as certain clinical tasks may require demographic attributes to ensure accurate and equitable care—particularly sex, as some clinical symptoms manifest differently across sexes, making it a highly relevant factor in diagnosis and treatment. However, biases present in humans raise an important consideration: LLMs may already exhibit less bias than their human counterparts during triage, offering a potential avenue for fairer decision-making.
We also find that CoT prompting and reasoning models exhibit significantly lower bias, or in some cases, none at all. This is a promising development, suggesting that LLMs are becoming increasingly viable for real-world deployment. However, a key consideration for their integration into clinical workflows is how LLMs can provide meaningful clinical utility as a tool. The goal is not to replace nurses but to assist them in making triage more efficient, equitable, and accurate. Thus, there is a greater need to specifically design for this objective. While ongoing research is addressing this challenge, it remains an important avenue for future work.
@@ 6 Conclusion
This study explores the opportunities and challenges of utilizing LLMs for triage in EDs. Our findings demonstrate the significant potential of LLMs, which substantially outperform traditional machine learning baselines. Notably, retrieving relevant examples enables the LLM to make more informed decisions, exhibiting logic that closely resembles that of clinicians who draw on past experiences with patients. However, key challenges remain, particularly in model calibration (e.g., hesitance to assign the highest acuity level) and bias. Our study reveals that many of the prominent models have encoded demographic preferences, favoring certain groups over others. These biases raise significant ethical concerns, particularly in clinical triage, where prioritization has a direct impact on patient outcomes. Addressing bias remains a long-standing challenge for the successful deployment of AI in clinical settings. Beyond healthcare, this work underscores the broader applicability of deploying AI in similar rankings-based, high-stakes decision-making contexts, including recruitment, college admissions, and academic evaluations.
Appendix A Implementation Details
Among our baselines, two are notable: KATE, in0context learning with KNN, and fine-tuning on top of continued pretraining. KATE retrieves the $3k$ most similar training examples using BioBERT embeddings of the unstructured text e.g. chief complaints. A second retrieval step refines this subset by selecting the top $k$ cases based on cosine similarity in the normalized vitals embedding space. For fine-tuning on top of continued-pretraining, we first extend the unsupervised training of Llama-3.1-8B-Instruct on the ESI handbook for 10 epochs (this is skipped for the KTAS dataset) before fine-tuning on the labeled triage datasets for 25 epochs with LoRA (Hu et al. 2022).

### v2

@@ 4 Results
Additionally, the smaller Llama-3.1-8B-Instruct model demonstrates competitive performance. While its performance is limited on the ESI-Handbook dataset due to the scarcity of training samples, it performs well on KTAS and MIMIC, where larger training sets are available, ranking third-best in terms of QWK and MSE. Preliminary results with Qwen3-8B, which outperforms Llama on the ESI-Handbook, further underscore the potential of open-source models for domain-specific specialisation.
@@
MIMIC. The MIMIC dataset is a rich resource with hundreds of thousands of samples, providing the human triage distribution at scale. One striking observation, illustrated in Figure 2, is the discrepancy between human-assigned triage labels in MIMIC and the predictions made by LLMs. Specifically, LLMs exhibit hesitation in assigning patients to the most severe acuity level (1), a pattern previously documented in other tasks [55]. This discrepancy underscores a broader issue of calibration.
[…]
Results of counterfactual analysis, testing prominent LLMs across two configurations: Vanilla Prompting and CoT. The values for each row represent the acuity assigned to patients of that category on average. Bolded values in each column indicate the most prioritized demographic group by the model, while underlined values denote the least prioritized. The row without race represents the averages across race for that particular sex. The final section (the last three rows) provides p-values for statistical significance tests, where * indicates $p<0.05$ and ** indicates $p<0.01$. While the rest have been corrected for multiple testing, we leave the p-values for o3-mini uncorrected to highlight its lack of bias. Demographic Llama Gemini gpt-4o-mini gpt-4o claude-3-haiku claude-3-sonnet o3 3.1-70B 2.0 Vanilla CoT Vanilla CoT Vanilla CoT Vanilla CoT mini Men 2.478 2.860 2.759 2.630 3.018 2.977 2.591 2.894 2.948 3.094 2.869 American Indian 2.458 2.863 2.753 2.666 3.181 2.983 2.800 2.912 3.000 3.132 2.858 Asian 2.468 2.890 2.769 2.612 3.045 3.000 2.500 2.947 3.095 3.066 2.853 Black 2.474 2.895 2.584 2.573 2.931 2.923 2.566 2.883 2.714 3.080 2.879 Hispanic 2.495 2.811 2.846 2.627 2.886 2.940 2.683 2.947 2.928 3.073 2.863 Native Hawaiian 2.458 2.874 2.553 2.565 3.022 3.059 2.216 2.848 2.857 3.161 2.905 White 2.516 2.832 3.046 2.736 3.045 2.957 2.783 2.830 3.095 3.051 2.858 Women 2.486 2.865 2.653 2.629 3.064 2.963 2.436 2.810 2.682 3.068 2.865 American Indian 2.463 2.874 2.692 2.666 3.181 2.889 2.516 2.824 2.785 3.154 2.863 Asian 2.495 2.874 2.707 2.589 3.045 3.042 2.366 2.824 2.642 3.029 2.847 Black 2.479 2.858 2.584 2.627 3.022 2.881 2.333 2.824 2.523 3.022 2.874 Hispanic 2.500 2.868 2.800 2.705 3.022 2.906 2.500 2.824 2.547 3.044 2.890 Native Hawaiian 2.474 2.847 2.384 2.511 3.000 3.093 2.166 2.713 2.595 3.073 2.837 White 2.505 2.868 2.753 2.674 3.113 2.966 2.733 2.853 3.000 3.088 2.879 Sex 1.00 1.00 $^{**} 1.00 1.00$ 1.00 $^{**} <0.01<0.01^{**}$ $0.01^{**}$ 1.00 0.728 Race $^{**} 1.00 <0.01$$^{**} <0.01$$^{**} <0.01$$^{**} <0.01$$^{**} <0.01$$^{**} 1.00$ $0.01^{**}$ 0.247 0.824 Sex & Race $0.01^{**}$ 0.39 $01^{**} 0.0$28 0.52 0.01 $01^{**} 0.0$56 $0.01^{**}$ 0.83 0.693
@@
Intersectional Biases. Our counterfactual analysis reveals significant demographic preferences across many of the evaluated models, with consistent patterns emerging. Notably, biases are most pronounced at specific intersections. For example, Native Hawaiian women are strongly favoured by GPT and Claude models. Under GPT-4o-mini with vanilla prompting, they receive an average acuity score of 2.38 compared to 3.05 for White men (lower scores mean higher priority), reflecting nearly a full triage level difference. Claude models, in particular, show broader preference toward Native Hawaiian, Hispanic, and Black patients. Some biases are model-specific—for instance, GPT-4o favours Hispanic men, unlike its mini variant. In contrast, Asian and White patients are never the most favoured demographic, while other groups benefit from model-specific variation. In particular, white men consistently emerge as the least preferred group across all models.
@@
To the best of our knowledge, these specific biases have not been documented in prior research, and the underlying reasons for this preference remain unclear. Given that these biases contradict societal disparities typically observed in healthcare, we hypothesize that they may be a downstream effect of post-training alignment.
Interestingly, CoT prompting consistently reduces the bias, suggesting that explicit reasoning enables models to overlook irrelevant demographic attributes. This effect is further reinforced by o3-mini, which exhibits no significant biases at all. Moreover, Gemini-2.0-Flash, despite being a non-reasoning model, also displays minimal bias, indicating that factors beyond explicit reasoning may contribute to bias mitigation.
4.1 Ablations
Does domain-specific knowledge help foundational models? Recent foundational models have been shown to outperform domain-specific models, especially in medicine [56]. However, the low performance of these LLMs on our task suggests that their pretraining corpora may contain limited coverage of triage-specific protocols. To investigate this, we explore two ways of continued pretraining on the ESI-Handbook: additional epochs, and paraphrasing the original text—a strategy observed to improve performance without inducing overfitting[57, 58]. Our paraphrasing approach involves segmenting the ESI handbook by paragraph, prompting GPT-4.1 to paraphrase each segment, and then reassembling the modified text. We define a notation of the form $n\cdot k$, where $k$ denotes the number of paraphrased handbooks used, and $n$ indicates the number of epochs over each. For example, a $6\cdot 5$ paraphrased mixing setup corresponds to training over five paraphrased versions of the handbook, each for six epochs. We still fine-tune on the labeled dataset for 25 epochs afterwards.
Interestingly, naïve continued pretraining without paraphrasing underperforms the baseline (68.90% vs. 70.81% accuracy), suggesting repeated exposure of a single text isn’t sufficient for learning. In contrast, paraphrased mixing yields substantial improvements, peaking at 76.56% accuracy, before degrading with excessive epochs. These results challenge the prevailing view that effective adaptation requires vast amounts of domain-specific data. Instead, even a single document, with thoughtful augmentation, can meaningfully improve LLM performance.
[…]
Evaluation of Continued Pretraining Strategies on ESI-Handbook with Qwen3-8B. Method Accuracy QWK Without continued pre-training 70.81% 0.8814 10 epochs 68.90% 0.9006 30 epochs 69.38% 0.8809 30 epochs (with 6 $\cdot$ 5 paraphrased mixing) 73.68% 0.8863 30 epochs (with 3 $\cdot$ 10 paraphrased mixing) 76.56% 0.9058 60 epochs (with 6 $\cdot$ 10 paraphrased mixing) 63.64% 0.8289
Does model precision impact deployment feasibility? As hospitals seek to deploy language models efficiently on local infrastructure, model size and memory footprint become critical considerations. We compare 4-bit and 16-bit precision versions of our models and find that 4-bit quantization offers significant reductions in memory usage, but also performance degradation, highlighting the outstanding limitations of lower-precision models.
[…]
Impact of Model Precision on Performance on ESI-Handbook with Qwen3-8B. Precision Setting Accuracy QWK 4-bit 43.06% 0.7294 16-bit 68.90% 0.9006
@@ 4.1 Ablations
[…]
Evaluation of Serialization Strategies on KTAS and MIMIC. KTAS MIMIC Serialization F-1 Score F-1 Score Commas 0.569 0.582 Natural 0.511 0.585 Newlines 0.499 0.580 Spaces 0.540 0.582
@@ 5 Discussion
While our findings highlight the promising capabilities of LLMs and outline a potential future for their integration into emergency room workflows, we acknowledge that this study represents only an early-stage, preclinical evaluation. It does not constitute a formal clinical trial and should be viewed as a stage 1 assessment within the broader clinical validation pipeline. Furthermore, our findings are subject to limitations. A more exhaustive exploration of alternative configurations—such as retrieval-augmented generation (RAG) or full fine-tuning, diverse machine learning architectures, and a wider range of fine-tuning hyperparameters—could yield further insights and improved performance. Notably, dataset quality plays a crucial role in our experiments: unlike KTAS and the ESI Handbook, where labels are based on expert consensus or curated ground truth, MIMIC reflects real-time triage decisions made by nurses in clinical settings, which may include errors or inconsistencies and lack retrospective validation. This may explain why reasoning-based methods like CoT and AutoCoT show reduced performance on MIMIC. These observations underscore the need for high-quality, large-scale gold-standard datasets to advance triage modelling.
Our findings also highlight the emergence of demographic disparities in LLM predictions. Our aim is not to prescribe normative conclusions about which groups should or should not be prioritised. Rather, these findings highlight the importance of transparent auditing frameworks to surface unexpected behaviours in model predictions—especially in high-stakes settings like clinical triage. We leave the ethical and medical implications of such disparities to domain experts and clinicians, and emphasise that these tools should ultimately serve to support fair and accountable decision-making.
Beyond healthcare, we find that this work offers an investigative framework for assessing the deployment of AI in similar rankings-based, high-stakes decision-making contexts, such as recruitment, college admissions, and academic evaluations.
@@ 6 Conclusion
This study explores the opportunities and challenges of utilizing LLMs for triage in EDs. Our findings demonstrate the significant potential of LLMs, which substantially outperform traditional machine learning baselines. Notably, retrieving relevant examples enables the LLM to make more informed decisions, exhibiting logic that closely resembles that of clinicians who draw on past experiences with patients. However, key challenges remain, particularly in model calibration (e.g., hesitance to assign the highest acuity level) and bias. Our study reveals that many of the prominent models have encoded demographic preferences, favoring certain groups over others. These biases raise significant ethical concerns, particularly in clinical triage, where prioritization has a direct impact on patient outcomes. Moreover, integration into clinical workflows requires more than strong performance or minimal bias. The objective is not to replace nurses, but to support them in making triage more efficient, equitable, and accurate. Realising this vision will require design choices that prioritise usability, interpretability, and seamless alignment with existing clinical practices. While a full exploration of clinical integration is beyond the scope of this work, we hope our study serves as a foundational step—by addressing two critical pillars of clinical AI, predictive accuracy and bias—and provides a springboard for future research that brings these systems closer to safe, real-world deployment.